# DeepPrune: Parallel Scaling without Inter-trace Redundancy

### v1

@@ Abstract
Parallel scaling has emerged as a powerful paradigm to enhance reasoning capabilities in large language models (LLMs) by generating multiple Chain-of-Thought (CoT) traces simultaneously. However, this approach introduces significant computational inefficiency due to inter-trace redundancy—our analysis reveals that over 80% of parallel reasoning traces yield identical final answers, representing substantial wasted computation. To address this critical efficiency bottleneck, we propose DeepPrune, a novel framework that enables efficient parallel scaling through dynamic pruning. Our method features a specialized judge model trained with focal loss and oversampling techniques to accurately predict answer equivalence from partial reasoning traces which realizes 0.87 AUROC on equivalence prediction, combined with an online greedy clustering algorithm that dynamically prunes redundant paths while preserving answer diversity. Comprehensive evaluations across three challenging benchmarks (AIME 2024, AIME 2025, and GPQA) and multiple reasoning models demonstrate that DeepPrune achieves remarkable token reduction by over 80% compared to conventional consensus sampling on most cases, while maintaining competitive accuracy within 3 percentage points. Our work establishes a new standard for efficient parallel reasoning, making high-performance reasoning more efficient. Our code and data are here: https://deepprune.github.io.
@@
Shangqing Tu1∗,Yaxuan Li2∗, Yushi Bai1, Lei Hou1, Juanzi Li1 1Tsinghua University, 2ShanghaiTech University https://deepprune.github.io thanks: Equal contribution.
@@ 1 Introduction
We conduct comprehensive experiments to prove the effectiveness of DeepPrune across diverse settings. In offline evaluation, our judge model achieves an average AUROC of 0.8701 and TNR of 0.8186 at FNR=0.2 when using reasoning words with focal loss and oversampling, significantly outperforming baseline methods. More importantly, in online reasoning tasks across three challenging benchmarks (AIME 2024, AIME 2025, and GPQA) and three state-of-the-art reasoning models (DeepSeek-8B, Qwen3-32B, and GPT-OSS-20B), DeepPrune reduces token consumption over 80% compared to cons@512 which samples 512 traces and conduct majority voting in most cases, while maintaining comparable accuracy (within 3 points). Notably, on AIME25 dataset, DeepPrune achieves up to 91.6% token reduction even with accuracy improvements, substantially outperforming strong baselines like DeepConf Fu et al. (2025b).
Our contributions are threefold: (1) We identify and quantify the pervasive problem of inter-trace redundancy in parallel reasoning, revealing that over 80% of computational resources are wasted on generating equivalent reasoning paths. (2) We propose DeepPrune, a novel framework that combines a trained judge model with online greedy clustering to efficiently prune redundant reasoning traces while preserving answer diversity. (3) Extensive offline and online experiments show that our method reduces token consumption by up to 95% without compromising accuracy, significantly outperforming existing baselines across multiple reasoning benchmarks and model architectures.
@@ 2 Related Work
Parallel Scaling. Parallel scaling has emerged as a pivotal paradigm for enhancing reasoning performance through concurrent generation of multiple reasoning traces Chen et al. (2024a); Pan et al. (2025); Zheng et al. (2025). Self-Consistency Wang et al. (2022) pioneered majority voting over diverse reasoning paths, while Best-of-$N$ sampling Brown et al. (2024) extended this concept through explicit candidate ranking. There are also tree-based exploration methods like ToT Yao et al. (2023).
Efficient Reasoning. Efficient reasoning methods Feng et al. (2025); Sui et al. (2025); Liu et al. (2025); Qu et al. (2025) aim to optimize the accuracy-compute trade-off during inference Kang et al. (2025); Srivastava et al. (2025); Li et al. (2025); Zhang et al. (2025); Wen et al. (2025). Prior research has explored reducing token usage in individual reasoning traces, such as through length-conscious fine-tuning Liu et al. (2024); Arora and Zanette (2025); Aggarwal and Welleck (2025); Xia et al. (2025) or training-free prompting techniques Renze and Guven (2024); Han et al. (2024); Xu et al. (2025); Fu et al. (2025a); Aytes et al. (2025). Another line of work improves parallel scaling efficiency by early-stopping redundant samples via confidence estimates Fu et al. (2025b); Yang et al. (2025b) or by refining aggregation strategies Wang et al. (2024, 2025b). While these methods address intra-trace verbosity or sample quantity reduction, they do not explicitly model redundancy between parallel reasoning paths. Our work directly targets this inter-trace redundancy, enabling proactive pruning while preserving answer diversity.
@@ 3.2 Inter-trace Redundancy: The Efficiency Bottleneck of Parallel Reasoning
Parallel Reasoning Trace Collection. To uncover the phenomenon of inter-trace redundancy, we conduct a reasoning trace collection process. We select four widely-used reasoning models: DeepSeek-R1-Distill-Llama-8B Guo et al. (2025), Qwen3-4B-Thinking-2507 Yang et al. (2025a), GLM-4.5-Air Zeng et al. (2025) and QwQ-32B. These models are evaluated on 115 problems sampled from reasoning benchmarks including MATH500 Hendrycks et al. (2021), AIME24, AIME25 AIME (2025), and GPQA Rein et al. (2024). For each problem, we generate 16 parallel reasoning traces per model. The traces for each problem are paired exhaustively, resulting in $\binom{16}{2}=120$ pairs per problem. Thus, for each model, we obtain around 13,000 pairs in total. To determine answer equivalence, we use the rule-based reward function from DeepScaleR Luo et al. (2025) to verify whether the final answers of each pair are identical. More details are in Appendix A.
The results, summarized in Figure LABEL:fig:subfig1, reveal a striking dominance of same-answer pairs across all models, with ratios exceeding 80% in most cases. Specifically, GLM-4.5-Air has 94.5% same-answer pairs. This severe class imbalance highlights that even with a modest number of samples (16 per problem), a large proportion of computational resources are wasted on generating redundant reasoning paths. The high prevalence of same-answer pairs (over 80% on average) underscores inter-trace redundancy as a critical efficiency bottleneck in parallel reasoning.
@@ 4.1.1 Trace Pair Data Collection
where $R(o_{i},o_{j})$ is a reward function that verifies answer equivalence using rule-based methods from DeepScaler Luo et al. (2025), and $o_{i},o_{j}$ are final answers derived from traces $t_{i}$ and $t_{j}$, respectively.
@@
Our training data is collected exclusively from DeepSeek-R1-Distill-Llama-8B Guo et al. (2025) outputs, while traces from other models are reserved for testing cross-model generalization.
@@ 4.2.1 Early Stopping via Greedy Clustering
where $t_{h}^{(j)}$ are randomly sampled top-$p$ traces from cluster $c_{j}$ with $p=\min(K_{1},|c_{j}|)$. If $\max_{j}\text{sim}(t_{i},c_{j})>\tau$, we assign $t_{i}$ to the most similar cluster $\arg\max_{j}\text{sim}(t_{i},c_{j})$; otherwise, we create a new cluster if the maximum number of clusters $K$ has not been reached. If $K$ is reached, we will terminate the clustering process. All the hyper-parameters are listed in Appendix C.
4.2.2 Majority Voting for Final Answer
After clustering, we need to select a final answer from the remaining traces. We observe two kinds of errors from our judge model: (1) Most pairs are classified as equivalent, so the largest cluster has too many traces. (2) All trace pairs are predicted as different, therefore each cluster only has one trace. To deal with these situations, we first select the largest cluster $c_{\text{max}}$ ($|c|$ means the number of reasoning traces in $c$):
$c_{\text{max}}=\arg\max_{c\in C}|c|$
To conduct voting without too many identical traces, we only let the top-$k^{*}$ traces in $c_{\text{max}}$ to finish reasoning, where $k^{*}=\min(|c_{\text{max}}|,K_{2})$. Besides, if all clusters are singletons, i.e. $|c|=1,\forall c\in C$, which means the judge model is highly likely wrong, we abandon the clustering results and just sample $k^{*}=K_{3}$ traces from $S$ for final reasoning. Finally, we apply majority voting on the final answers of those finished traces:
$o_{\text{final}}=\text{MajorityVote}(\{o_{1},o_{2},\dots,o_{k^{*}}\})$
This approach ensures that we can invest computational resources primarily in promising reasoning paths even if the judge model may produce wrong prediction, which reduces token consumption in parallel reasoning while preserving correctness.
@@ 4.2.2 Majority Voting for Final Answer
Table 1: The offline evaluation results of the judge model across different truncation methods and training strategies, which reports the average AUROC and TNR@0.2 metrics for three reasoning models using two truncation types: first-500 tokens and first-25 reasoning words, with the combination of focal loss and oversampling.
@@ 5.1 Experimental Setup
Settings. Our evaluation includes both offline assessment of the judge model’s predictive performance and online testing of the full pruning framework during reasoning tasks. For offline evaluation, we train our judge model on reasoning traces from DeepSeek-R1-Distill-Llama-8B with datasets collected in Section 4.1.1 and evaluate its generalization capability on three distinct reasoning models that are not included in the training data: Qwen3-4B-Thinking-2507, QwQ-32B, and GLM-4.5-Air. For online reasoning experiments, we evaluate three reasoning models: DeepSeek-8B [^1] https://huggingface.co/deepseek-ai/DeepSeek-R1-0528-Qwen3-8B Guo et al. (2025), Qwen3-32B Yang et al. (2025a), and GPT-OSS-20B Agarwal et al. (2025) across the three benchmarks: AIME 2024 MAA (2024), AIME 2025 AIME (2025), and GPQA Rein et al. (2024). We generate 512 parallel reasoning traces per problem for baseline methods and apply DeepPrune with a redundancy threshold $\tau=0.5$.
@@
$\Delta\text{Token}\%=\frac{\text{Tokens}_{\text{pruned}}-\text{Tokens}_{\text{origin}}}{\text{Tokens}_{\text{origin}}}\times 100\%$
Baselines. We compare DeepPrune against several competitive baselines: (1) Sampling Methods: cons@512 which samples 512 parallel traces with majority voting for self-consistency Wang et al. (2022), serves as our primary baseline for token reduction calculations. (2) Confidence-based Pruning Methods: DeepConf-high and DeepConf-low are confidence-based early stopping methods with high or low threshold for pruning. These baselines represent the state-of-the-art in efficient reasoning methods. We ensure fair comparison by using identical model checkpoints and configurations with the DeepConf paper with details in Appendix B.
@@
   Token ($\times 10^{8}$) 0.42 0.35 2.54 0.26 0.23 1.00 0.42 0.38 2.20
   $\Delta\text{Token}\%$ -88.3% -91.6% -76.7% -86.4% -91.4% -85.6% -79.6% -82.2% -52.5%
   Accuracy 86.7% 83.3% 63.1% 90.0% 90.0% 70.2% 90.0% 93.3% 68.7%
@@ 5.2 Offline Experiment Results
As shown in Table 1, we have three observations: (1) Our best configuration, which uses first-25 reasoning words with focal loss and oversampling, achieves superior performance with an average AUROC of 0.8701 and TNR@0.2 of 0.8186 across all models. This represents a substantial improvement over the preliminary zero-shot LLM judgment (AUROC=0.66) reported in Figure LABEL:fig:subfig3, validating the necessity of specialized training for redundancy prediction. (2) The comparison between first-500 tokens and first-25 reasoning words reveals a clear advantage for reasoning-aligned truncation. This suggests that structural alignment with reasoning steps provides more reliable signals for predicting answer equivalence compared to fixed-length token windows. (3) The ablation study on training strategies demonstrates the critical importance of addressing class imbalance. The combination of focal loss and oversampling consistently delivers the best performance across both truncation types and all reasoning models. Notably, using oversampling alone significantly degrades performance (AUROC drops to 0.7610 for tokens and 0.7983 for reasoning words), indicating that simply balancing the dataset distribution is insufficient without proper loss weighting. Focal loss alone provides moderate improvements, but the synergistic combination with oversampling yields the most robust results.
$\tau$ Qwen3-32B on AIME24 Qwen3-32B on AIME25
Greedy Clustering w/ Majority Voting Greedy Clustering w/ Majority Voting
@@ 5.3 Online Experiment Results
Table 2 presents the online evaluation results, from which we can get several key findings.
Substantial Token Reduction with Minimal Accuracy Loss. DeepPrune achieves remarkable token savings while maintaining competitive accuracy across all experimental settings. Specifically, DeepPrune reduces token consumption by over 80% in most cases compared to the cons@512 sampling baseline. The most significant reductions are observed on AIME datasets, where DeepPrune achieves 79.6%-91.6% token savings with negligible accuracy drop (within 3 percentage points). For instance, on Qwen3-32B with AIME25, DeepPrune reduces tokens by 91.4% while even improving accuracy from 80.0% to 90.0%. This demonstrates that our method effectively identifies and eliminates redundant reasoning paths without compromising solution quality.
@@
Cross-Model Generalization. The consistent performance across three distinct reasoning models (DeepSeek-8B, Qwen3-32B, and GPT-OSS-20B) validates the generalizability of our approach. Particularly noteworthy is that the judge model of DeepPrune was trained purely on the reasoning traces of DeepSeek-R1-Distill-Llama-8B, which means all the tested models are out-of-distribution, providing a robust solution for efficient parallel reasoning.

### v2

@@ Abstract
Parallel scaling has emerged as a powerful paradigm to enhance reasoning capabilities in large language models (LLMs) by generating multiple Chain-of-Thought (CoT) traces simultaneously. However, this approach introduces significant computational inefficiency due to inter-trace redundancy—our analysis reveals that over 80% of parallel reasoning traces yield identical final answers, representing substantial wasted computation. To address this critical efficiency bottleneck, we propose DeepPrune, a novel framework that enables efficient parallel scaling through dynamic pruning. Our method features a specialized judge model trained with out-of-distribution data (AIME 2022, AIME 2023, and MATH 500) using oversampling techniques to accurately predict answer equivalence from partial reasoning traces, achieving 0.7072 AUROC on unseen reasoning models. Combined with an online greedy clustering algorithm that dynamically prunes redundant paths while preserving answer diversity. Comprehensive evaluations across three challenging benchmarks (AIME 2024, AIME 2025, and GPQA) and multiple reasoning models demonstrate that DeepPrune achieves remarkable token reduction of 65.73%–88.50% compared to conventional consensus sampling, while maintaining competitive accuracy within 3 percentage points. Our work establishes a new standard for efficient parallel reasoning, making high-performance reasoning more efficient. Our code and data are here: https://deepprune.github.io/.
@@
Shangqing Tu1thanks:   Equal Contribution.,Yaxuan Li2∗, Yushi Bai1, Lei Hou1thanks:   Corresponding author., Juanzi Li1 1Tsinghua University, 2ShanghaiTech University {tsq25,bys22}@mails.tsinghua.edu.cn liyx12023@shanghaitech.edu.cn,{houlei,lijuanzi}@tsinghua.edu.cn https://deepprune.github.io
@@ 1 Introduction
We conduct comprehensive experiments to prove the effectiveness of DeepPrune across diverse settings. To ensure robust generalization, we train our judge model on fully out-of-distribution data (AIME 2022, AIME 2023, and MATH 500) and evaluate it on unseen reasoning models and benchmarks. In offline evaluation, our judge model achieves an average AUROC of 0.7072 and TPR of 0.5063 when using first-800 tokens with oversampling, demonstrating strong cross-model generalization to three unseen reasoning models. More importantly, in online reasoning tasks across three challenging benchmarks (AIME 2024, AIME 2025, and GPQA) and three state-of-the-art reasoning models (DeepSeek-8B, Qwen3-32B, and GPT-OSS-20B), DeepPrune reduces token consumption by 65.73% to 88.50% compared to cons@512 which samples 512 traces and conducts majority voting, while maintaining comparable accuracy (within 3.4 percentage points). Notably, on AIME24 and AIME25 datasets, DeepPrune achieves up to 88.5% token reduction while maintaining or improving accuracy, substantially outperforming strong baselines like DeepConf Fu et al. (2025b).
Our contributions are threefold: (1) We identify and quantify the pervasive problem of inter-trace redundancy in parallel reasoning, revealing that over 80% of computational resources are wasted on generating equivalent reasoning paths. (2) We propose DeepPrune, a novel framework that combines a trained judge model with online greedy clustering to efficiently prune redundant reasoning traces while preserving answer diversity. (3) Extensive offline and online experiments show that our method reduces token consumption by up to 88.5% without compromising accuracy, significantly outperforming existing baselines across multiple reasoning benchmarks and model architectures.
@@ 2 Related Work
Parallel Scaling. Parallel scaling has emerged as a pivotal paradigm for enhancing reasoning performance through concurrent generation of multiple reasoning traces Chen et al. (2024a); Pan et al. (2025); Zheng et al. (2025). Self-Consistency Wang et al. (2022) pioneered majority voting over diverse reasoning paths, while Best-of-$N$ sampling Brown et al. (2024) extended this concept through explicit candidate ranking. There are also tree-based exploration methods like ToT Yao et al. (2023) that dynamically branch reasoning paths.
Efficient Reasoning. Efficient reasoning methods Feng et al. (2025); Sui et al. (2025); Liu et al. (2025); Qu et al. (2025) aim to optimize the accuracy-compute trade-off during inference Kang et al. (2025); Srivastava et al. (2025); Li et al. (2025). Prior research has explored reducing token usage in individual reasoning traces, such as through length-conscious fine-tuning Liu et al. (2024); Arora and Zanette (2025); Aggarwal and Welleck (2025); Xia et al. (2025) or training-free prompting techniques Renze and Guven (2024); Han et al. (2024); Xu et al. (2025); Fu et al. (2025a); Aytes et al. (2025). Another line of work improves parallel scaling efficiency by early-stopping redundant samples via confidence estimates Fu et al. (2025b); Yang et al. (2025b) or by refining aggregation strategies Wang et al. (2024, 2025b). While these methods address intra-trace verbosity or sample quantity reduction, they do not explicitly model redundancy between parallel reasoning paths. Our work directly targets this inter-trace redundancy, enabling proactive pruning while preserving answer diversity.
@@ 3.2 Inter-trace Redundancy: The Efficiency Bottleneck of Parallel Reasoning
Parallel Reasoning Trace Collection. To uncover the phenomenon of inter-trace redundancy, we conduct a reasoning trace collection process. We select four widely-used reasoning models: Deepseek-R1-Distill-Llama-8B, Qwen3-4B-Thinking-2507, GLM-4.5-Air and QWQ-32B . These models are evaluated on over 100 problems from reasoning benchmarks including Math500 Hendrycks et al. (2021), AIME24, AIME25 AIME (2025), and GPQA Rein et al. (2024). For each problem, we generate 16 parallel reasoning traces per model. The traces for each problem are paired exhaustively, resulting in $\binom{16}{2}=120$ pairs per problem. Thus, for each model, we obtain around 12,000 pairs in total. To determine answer equivalence, we use the rule-based reward function from DeepScaleR Luo et al. (2025) to verify whether the final answers of each pair are identical. More details are in Appendix A.
The results, summarized in Figure LABEL:fig:subfig1, reveal a striking dominance of same-answer pairs across all models, with ratios exceeding 80% in most cases. Specifically, GLM-4.5-Air Zeng et al. (2025) has 94.5% same-answer pairs. This severe class imbalance highlights that even with a modest number of samples (16 per problem), a large proportion of computational resources are wasted on generating redundant reasoning paths. The high prevalence of same-answer pairs (over 80% on average) underscores inter-trace redundancy as a critical efficiency bottleneck in parallel reasoning.
@@ 4.1.1 Trace Pair Data Collection
where $R(o_{i},o_{j})$ is a reward function that verifies answer equivalence using rule-based methods from DeepScaler, and $o_{i},o_{j}$ are the final answers derived from traces $t_{i}$ and $t_{j}$ respectively.
@@
Our training data is collected exclusively from Deepseek-R1-Distill-Llama-8B outputs using out-of-distribution datasets (AIME 2022, AIME 2023, and MATH 500), while traces from other models and different datasets are reserved for testing cross-model generalization.
@@ 4.2.1 Early Stopping via Greedy Clustering
where $t_{h}^{(j)}$ are randomly sampled top-$p$ traces from cluster $c_{j}$ with $p=\min(20,|c_{j}|)$. If $\max_{j}\text{sim}(t_{i},c_{j})>\tau$, we assign $t_{i}$ to the most similar cluster $\arg\max_{j}\text{sim}(t_{i},c_{j})$; otherwise, we create a new cluster if the maximum number of clusters $K$ has not been reached. If $K$ is reached, we will terminate the clustering process.
@@
Table 1: The offline evaluation results of the judge model across different truncation methods and training strategies, which reports the average AUROC and TNR@0.2 metrics for three reasoning models using two truncation types: top-500 tokens and top-25 reasoning words, with the combination of focal loss and oversampling.
4.2.2 Majority Voting for Final Answer
After clustering, we need to select a final answer from the remaining traces. Since our judge model may give wrong prediction, we observe two kinds of errors: (1) Most pairs are classified as equivalent, so the largest cluster has too many traces. (2) All trace pairs are predicted as different, therefore each cluster only has one trace. To deal with these situations, we first select the largest cluster $c_{\text{max}}$ ($|c|$ means the number of reasoning traces in $c$):
$c_{\text{max}}=\arg\max_{c\in C}|c|$
To conduct voting without too many identical traces, we only let top-$q_{1}$ traces in $c_{\text{max}}$ to finish reasoning, where $q_{1}=\min(|c_{\text{max}}|,20)$. Besides, if all clusters are singletons, i.e. $|c|=1,\forall c\in C$, which means the judge model is highly likely wrong, we just sample $q_{2}=32$ traces from $S$ for final reasoning. Finally, we apply majority voting on the final answers of those finished traces:
$o_{\text{final}}=\text{MajorityVote}(\{o_{1},o_{2},\dots,o_{q}\})$
where $q=q_{1}$ or $q=q_{2}$ depending on situations.
This approach ensures that we can invest computational resources primarily in promising reasoning paths even if the judge model may produce wrong prediction, which reduces token consumption in parallel reasoning while preserving answer quality.
@@ 5.1 Experimental Setup
Settings. Our evaluation includes both offline assessment of the judge model’s predictive performance and online testing of the full pruning framework during reasoning tasks. For offline evaluation, we train our judge model on reasoning traces from Deepseek-R1-Distill-Llama-8B and evaluate its generalization capability on three distinct reasoning models: Qwen3-4B-Thinking-2507, QWQ-32B, and glm-4.5-air. We use reasoning traces collected from challenging reasoning benchmarks comprising over 1,000 problems. For online reasoning experiments, we evaluate three reasoning models: DeepSeek-8B Guo et al. (2025), Qwen3-32B Yang et al. (2025a), and GPT-OSS-20B Agarwal et al. (2025) across the three benchmarks: AIME 2024 MAA (2024), AIME 2025 AIME (2025), and GPQA Rein et al. (2024). We generate 512 parallel reasoning traces per problem for baseline methods and apply DeepPrune with a redundancy threshold $\tau=0.5$.
@@
$\Delta\text{Token}\%=\frac{\text{Tokens}_{\text{new}}-\text{Tokens}_{\text{origin}}}{\text{Tokens}_{\text{origin}}}\times 100\%$
Baselines. We compare DeepPrune against several competitive baselines: (1) Sampling Methods: cons@512 which samples 512 parallel traces with majority voting for self-consistency Wang et al. (2022), serves as our primary baseline for token reduction calculations. (2) Confidence-based Pruning Methods: DeepConf-high and DeepConf-low are confidence-based early stopping methods with high or low threshold for pruning. These baselines represent the state-of-the-art in efficient reasoning methods. We ensure fair comparison by using identical model checkpoints and experimental configurations with the DeepConf Fu et al. (2025b).
@@
   Token ($\times 10^{8}$) 0.42 0.35 2.54 0.26 0.23 - 0.42 0.38 2.20
   $\Delta\text{Token}\%$ -88.3% -91.6% -76.7% -86.4% -91.4% -% -79.6% -82.2% -52.5%
   Accuracy 86.7% 83.3% 63.1% 90.0% 90.0% -% 90.0% 93.3% 68.7%
@@ 5.2 Offline Experiment Results
As shown in Table 1, we have three observations: (1) Our best configuration, which uses top-25 reasoning words with focal loss and oversampling, achieves superior performance with an average AUROC of 0.8701 and TNR@0.2 of 0.8186 across all models. This represents a substantial improvement over the preliminary zero-shot LLM judgment (AUROC=0.66) reported in Figure LABEL:fig:subfig3, validating the necessity of specialized training for redundancy prediction. (2) The comparison between top-500 tokens and top-25 reasoning words reveals a clear advantage for reasoning-aligned truncation. This suggests that structural alignment with reasoning steps provides more reliable signals for predicting answer equivalence compared to fixed-length token windows. (3) The ablation study on training strategies demonstrates the critical importance of addressing class imbalance. The combination of focal loss and oversampling consistently delivers the best performance across both truncation types and all reasoning models. Notably, using oversampling alone significantly degrades performance (AUROC drops to 0.7610 for tokens and 0.7983 for reasoning words), indicating that simply balancing the dataset distribution is insufficient without proper loss weighting. Focal loss alone provides moderate improvements, but the synergistic combination with oversampling yields the most robust results.
Threshold Qwen3-32B on AIME24 Qwen3-32B on AIME25
Greedy Clustering w. Majority Voting Greedy Clustering w. Majority Voting
@@ 5.3 Online Experiment Results
Table 2 presents the online evaluation results, from which we can get several key findings:
Substantial Token Reduction with Minimal Accuracy Loss. DeepPrune achieves remarkable token savings while maintaining competitive accuracy across all experimental settings. Specifically, DeepPrune reduces token consumption by 52.5% to 91.6% compared to the cons@512 sampling baseline. The most significant reductions are observed on AIME datasets, where DeepPrune achieves 79.6%-91.6% token savings with negligible accuracy drop (within 3 percentage points). For instance, on Qwen3-32B with AIME25, DeepPrune reduces tokens by 91.4% while even improving accuracy from 80.0% to 90.0%. This demonstrates that our method effectively identifies and eliminates redundant reasoning paths without compromising solution quality.
@@
Cross-Model and Cross-Dataset Generalization. The consistent performance across three distinct reasoning models (DeepSeek-8B, Qwen3-32B, and GPT-OSS-20B) and three diverse benchmarks (AIME24, AIME25, GPQA) validates the generalizability of our approach. Particularly noteworthy is that the judge model of DeepPrune was trained purely on the reasoning traces of Deepseek-R1-Distill-Llama-8B, which means all the tests are out-of-distribution, providing a practical solution for efficient parallel reasoning